OpenStudy (anonymous):

Tan[(1/2)sin^-1(4/5)]

OpenStudy (anonymous):

@terenzreignz

terenzreignz (terenzreignz):

Ok this time, recall the half-angle identities...

terenzreignz (terenzreignz):

Maybe it would be to your advantage to use one of these \[\huge \tan\left[\frac12 x\right]=\frac{1-\cos(x)}{\sin(x)}=\csc(x) -\cot(x)\] Whichever of these may be more useful for you.

OpenStudy (anonymous):

the second one is familiar

terenzreignz (terenzreignz):

\[\large \tan\left[ \frac12 \sin^{-1}\left(\frac45\right)\right]=\frac{1-\cos\left[ \sin^{-1}\left(\frac45\right)\right]}{\sin\left[ \sin^{-1}\left(\frac45\right)\right]}=\csc\left[ \sin^{-1}\left(\frac45\right)\right] -\cot\left[ \sin^{-1}\left(\frac45\right)\right]\]

terenzreignz (terenzreignz):

Never mind the cut-off part, the first one is easier, anyway. Now what is \[\huge \sin\left[ \sin^{-1}\left(\frac45\right)\right]\]?

OpenStudy (anonymous):

4/5

terenzreignz (terenzreignz):

That's right. Now what is \[\huge \cos\left[ \sin^{-1}\left(\frac45\right)\right]\]?

OpenStudy (anonymous):

3/5

terenzreignz (terenzreignz):

oh... I was about to draw a triangle, but it seems you have it down. Now all you have to do is plug in... \[\large \tan\left[ \frac12 \sin^{-1}\left(\frac45\right)\right]=\frac{1-\cos\left[ \sin^{-1}\left(\frac45\right)\right]}{\sin\left[ \sin^{-1}\left(\frac45\right)\right]}\]

OpenStudy (anonymous):

wow. Ok I understand

OpenStudy (anonymous):

very simple, once you get it. Thanks a lot!

terenzreignz (terenzreignz):

No problem :)